# An Integrated Deep Learning Model for Skin Cancer Detection Using Hybrid Feature Fusion Technique

### v1

@@ 1 Introduction
Cancer happens to be one of the most lethal diseases that may occur in humans. It is the world’s second leading cause of death 1. There are many different kinds of cancer in the human body, and skin cancer is one of the most frequent, lethal, and rapidly spreading malignancies that can lead to death. Skin cancer develops from abnormal growth of cells that make up skin, which shields the body from heat, light and infection 2. It is caused by factors such as genetic factors, alcohol consumption, smoking, infections, allergies, viruses, environmental change, physical exercise, exposure to ultraviolet light, and so on. Fungal growth and bacteria are the most likely reasons. In addition, Skin cancer is also caused by unusual swellings of the human body 3.
@@
There are various kinds of skin cancer, and among them, melanoma is the most widespread and deadliest type of skin cancer. Melanoma evolves from maliciously transformed skin melanocytes. Melanoma cells are known to spread to other areas of the body, including the liver, lungs, spleen, or brain 2, 4. Melanocytes are responsible for the production of dark pigments on the hair, skin, eyes, and different areas of the body. Melanoma lesions tend to be black or brown in color. But in a few instances, it is also found in different colors such as azure, royal purple, rosy pink, or even colorless 5.
@@
According to the World Health Organization (WHO), cancer is one of the leading causes of death worldwide and is responsible for about 10 million deaths annually 6. It is responsible for about 1 in every 6 deaths worldwide. Between 2008 and 2030, global cancer-related deaths are expected to go up by 45% 7. According to the WHO, one out of every three people in the globe has skin cancer. Melanoma and non-melanoma are the most typical cancers about 300,000 new cases were found in 2018. Melanoma cost the lives of an estimated 2490 females and 4740 males in 20195.
@@
Deep learning (DL) is the method by which a computer model directly learns to perform categorization tasks from images, text, or sound 8, 9, 10, 11. Multi-layered neural network architectures and labeled data are used to train models. DL lessens the requirement for some of the data preparation that machine learning generally demands 12, 13. Unstructured data like text and images can be processed by it. Gradient descent and backpropagation are used to update and tune the DL algorithm for accuracy, enabling it to make predictions with greater precision. DL capabilities have had a significant positive impact on the healthcare sector. It has several uses in various areas of medical imaging because of its high level of precision. By using these techniques, the diagnosing procedure is expedited and human error is decreased.
@@
The most cutting-edge methods are without a doubt DL algorithms, which have lately achieved remarkable classification and segmentation results not only for medical images but also for others applications 14, 15. So, the primary reason to create the proposed hybrid DL system is to use it as a diagnostic instrument to assist individuals who make decisions in health and medical facilities in quickly and accurately identifying melanoma in dermatological images. This reduces the burden on pathologists and hospitals at a time when the number of individuals infected with skin cancer is increasing dramatically. The following are the primary contributions of this study:
@@ 2 Literature Review
In order to detect skin cancer, Yuvika Gautam et al. 16 presented FusionEXNet, a novel fused deep-learning model. In order to take use of their capacities for feature extraction from dermoscopic images, the model combines two architectures: XceptionNet and EfficientNetV2S. Approximately 10,000 high-resolution photos from seven different kinds of skin lesions are included in the HAM10000 dataset, which was used for training and evaluation. SmoothGrad and Faster Score-CAM, two Explainable Artificial Intelligence (XAI) approaches, were integrated into the model to improve interpretability and provide insights into the model’s decision-making process. FusionEXNet demonstrated superior performance in accuracy, achieving 90.83%, compared to XceptionNet (88.82%) and EfficientNetV2S (88.01%).
@@
In order to overcome the problems of light reflections, colour fluctuations, and lesion diversity, Md Shahin Ali et al. 3 suggested a deep convolutional neural network (DCNN) model for identifying benign and malignant skin lesions from dermoscopic pictures. Using filters, preprocessing eliminates noise and artefacts from the input images, normalises them, and applies data augmentation to grow the dataset size and enhance classification accuracy. The model’s performance is contrasted with that of various other transfer learning models, such as MobileNet, AlexNet, ResNet, VGG-16, and DenseNet. The suggested DCNN model outperformed the other models in terms of robustness and reliability using the HAM10000 dataset, with testing accuracy of 91.93% and training accuracy of 93.16%.
@@
Rehan Ashraf et al. 2 suggested an intelligent Region of Interest (ROI)-based melanoma detection system to overcome the difficulties caused by nevi’s difficult-to-see visual appearance and deep learning models’ restricted data availability. The methodology ensures that only pertinent melanoma features are incorporated for system training by using an updated k-means algorithm to extract ROIs from dermoscopic pictures. This ROI-based method concentrates on the areas that include melanoma cells, which improves feature discrimination. Subsequently, a Convolutional Neural Network (CNN) transfer learning model is trained on the ROI pictures from the DermIS and DermQuest datasets using data augmentation approaches. The suggested system outperforms current techniques that depend on comprehensive picture categorisation in terms of classification accuracy.
@@
In this study 17, dermoscopy images from the HAM10000 dataset were used to create a proprietary Convolutional Neural Network (CNN) model for skin cancer classification. The seven categories bearing and malignant that the CNN model was built to identify from the photos are essential for early detection and lowering the death rate from skin cancer. An Enhanced Super Resolution Generative Adversarial Network (ESRGAN) was used to preprocess the dataset in order to improve image quality by increasing the resolution of smaller sized images.
@@
In this study 18, a novel hybrid dynamic Bayesian Deep Learning model is introduced, utilizing Three-Way Decision (TWD) theory to adaptively apply different UQ methods and neural networks in distinct classification phases. The model’s effectiveness is validated on two skin cancer datasets, achieving: accuracy of 88.95% and F1-score of 89.00% for first dataset and accuracy of 90.96% and F1-score of 91.00% for second dataset.
@@
This study 19 uses dermoscopic images to offer a two-stream deep neural network information fusion framework for multiclass skin cancer classification. A pretrained DenseNet201 model is fed with the improved image quality through a fusion based contrast enhancement technique in the first stream of the methodology, which extracts features. An approach called skewness controlled mothflame optimisation is used to optimise these properties. A refined MobileNetV2 model’s features are taken out and further downsampled in the second stream through the use of a feature selection framework. A new parallel multimax coefficient correlation technique is employed to fuse discriminant characteristics from both networks. To classify lesion images, a multiclass extreme learning machine classifier is used. Three unbalanced skin datasets HAM10000, ISBI2018, and ISIC2019, respectively are used to evaluate the concept.
@@
This article 7 introduces two novel hybrid CNN models that incorporate a support vector machine (SVM) classifier at the output layer. The approach involves: Extracting features from both CNN models. Concatenating these features and inputting them into the SVM for final classification.The proposed models outperformed existing state-of-the-art CNN models achieving accuracies of 88.02% and 87.43%.
@@
This study 20 compared the diagnostic ability of dermatologists with a Faster Region-based Convolutional Neural Network (FRCNN) trained to categorise clinical photos of pigmented skin lesions. The dataset included 3551 patients’ 5846 photos, with lesions classified into six classes: benign hematoma/hemangioma, nevus, seborrhoeic keratosis, senile lentigo, and malignant melanoma. A test set including 666 photos was generated, and the remaining 4732 images were annotated with bounding boxes to form the training set. In a six-class classification, the FRCNN model outperformed doctors with board certification (79.5%) and dermatology trainees (75.1%), with an accuracy of 86.2%. FRCNN achieved 91.5% accuracy in binary classification (benign vs. malignant), with 83.3% sensitivity and 94.5% specificity, in that order. In comparison to dermatologists, it also showed a lower false positive rate of 5.5%. The results show that the FRCNN model performed better in picture classification tasks than 20 dermatologists, indicating that the public may benefit from using it in the future to improve skin cancer prognosis.
@@
The goal of this work 21 is to create deep learning models for the classification of dermal cell pictures in order to detect skin cancer. Deep learning techniques are used at the heart of the system to improve prediction accuracy through the use of a model-driven architecture hosted in the cloud. Building these models and using them to categorise dermal cells constitute the process; standard datasets are used for testing. With this method, practitioners may quickly develop and use deep learning models for early skin cancer identification.
@@
This work 4 addressed the difficulties caused by the fine-grained heterogeneity in diagnostic categories by developing a preprocessing pipeline for multiclass skin cancer classification. Image hair removal, dataset enhancement, and image resizing to accommodate various models were all part of the pipeline. To fine-tune the EfficientNet B0-B7 models for classification, transfer learning was used on pre-trained ImageNet weights using the HAM10000 dataset. With an F1 Score of 87% and Top-1 Accuracy of 87.91%, EfficientNet B4 had the best results when the models’ performance was assessed using Precision, Recall, Accuracy, F1 Score, and Confusion Matrices. More complicated models do not always perform better; models with intermediate complexity, such EfficientNet B4 and B5, showed notable superiority over more complex models.
@@
This research 22 utilizes images from various publicly available ISIC datasets to create a balanced dataset consisting of 10,500 images for training and testing. An ensemble of four convolutional neural network (CNN) architectures-ResNet50, EfficientNet B6, InceptionV3, and Xception was employed for melanoma classification.The experimental results demonstrate that the proposed ensemble model effectively classifies melanoma skin cancer with high accuracy, outperforming many state-of-the-art methods.
@@ 3.1 Dataset
Our dataset basically consists of two distinct categories of skin carcinoma images. DL requires a significant amount of data in order to produce a good result. The collection of skin cancer pictures, on the other hand, is critical. To address these issues, we used a freely available dataset of 3297 dermoscopic images from Kaggle 23. It includes 1800 images of benign and 1497 images of malignant categorized moles. The data was obtained via ISIC Archive. Each picture is 224x244 pixels in size. The professional pathologist confirmed the database’s ground truth.
@@ 3.2 Image Preprocessing
The main goal of the image preprocessing process is to make the data suitable for deep learning models and improve the initial medical images by getting rid of noise, air bubbles, and artifacts caused by gel that was sprayed before the picture was taken. To attain a high classifying rate, we removed artifacts and noise from the pictures. The elimination of this noise and artefacts was crucial to guaranteeing high-quality input images since they could mask important features required for classification. The image was smoothed using noise reduction algorithms, which also removed any random deviations that could confound the model. We employed feature scaling and turned the categorical label data into number data15. In order to ensure that the input data had consistent numerical ranges, we used feature scaling to normalise the pixel intensity values, which can vary significantly in photos. This step prevents gradient instability problems and helps models converge more quickly during training.
@@ 3.2.1 Class Labeling
We used label encoding techniques to transform the categorical data into numerical format because the labels in medical picture datasets are typically categorical (e.g., different forms of skin cancer). In order to enable the deep learning model to process and learn from the labels during training, each distinct category was given a numerical value. The labeling data collection is presented as follows: the image’s label 0 indicates benign patients, while 1 symbolizes malignant patients24.
@@ 3.2.2 Feature scaling
The feature scaling method is well recognized in the field of machine learning and pattern recognition for being used to normalize data. To avoid outliers, this approach sets all data items to the same scale and thus improves prediction quality. As the features in cancer datasets have a high variance, One of the pre-processing strategies for normalization is feature scaling. We divided the grayscale value of an image by 255 to standardize our image pixels between 0 and 1. As a result, the numbers will be relatively little, and the computation will be simpler and faster25, 26.
@@ 3.3 Transfer Learning
Transfer Learning has emerged as a popular method in Deep Learning (DL) that utilizes a pre-trained model for a similar task to train a new model. This approach is particularly useful in medical imaging, where training on large datasets such as ImageNet with all neural network parameters can be computationally intensive 6. The ImageNet dataset, for instance, contains 14 million images with 20,000 categories for visual recognition tasks 27.
In this study, we propose a model for skin lesion classification, which was trained and evaluated using advanced Convolutional Neural Networks (CNNs) such as Inception V3, Densenet121, and a Hybrid model 21, 3. To optimize our model, we replaced the last softmax layer with a sigmoid activation, reducing the classification output to two classes for identifying benign and malignant skin lesions.
The Inception V3 CNN model achieved remarkable accuracy of over 78.1% on the ImageNet dataset and has been widely used for image classification tasks 25. Its architecture includes inception modules that use filters of different sizes on the same input, and auxiliary classifiers for regularization weight loss ratio. The convolution layer implements factorization to minimize dimensionality and avoid overfitting.
Another CNN architecture utilized in our model is DenseNet121, which resolves the vanishing gradient problem by directly connecting every layer to every other layer. Instead of summing feature maps from previous layers, DenseNet121 concatenates them, resulting in a compact model with fewer parameters and enabling feature reuse 28.
@@ 3.4.1 Inception V3
InceptionV3 is a CNN-based DL model used for image classification. It is a highly renowned image detection model that has been developed over time by a number of researchers. On the ImageNet dataset, the model achieved more than 78.1% accuracy. Google’s InceptionV3 is the third version in the DL Convolutional Architecture series. The model was developed based on 1000 classes from the ImageNet dataset, which was trained on approximately 1 million images. It is made up of inception modules that apply filters of various sizes to the same amount of input. In the Inception network, the auxiliary classifiers add to the regularization weight loss ratio. Factorization was implemented in the convolution layer in order to minimize dimensionality and therefore overfitting 25.

### v2

@@ 1 Introduction
Cancer happens to be one of the most lethal diseases that may occur in humans. It is the world’s second leading cause of death [1]. There are many different kinds of cancer in the human body, and skin cancer is one of the most frequent, lethal, and rapidly spreading malignancies that can lead to death. Skin cancer develops from abnormal growth of cells that make up skin, which shields the body from heat, light and infection [2]. It is caused by factors such as genetic factors, alcohol consumption, smoking, infections, allergies, viruses, environmental change, physical exercise, exposure to ultraviolet light, and so on. Fungal growth and bacteria are the most likely reasons. In addition, Skin cancer is also caused by unusual swellings of the human body [3].
@@
There are various kinds of skin cancer, and among them, melanoma is the most widespread and deadliest type of skin cancer. Melanoma evolves from maliciously transformed skin melanocytes. Melanoma cells are known to spread to other areas of the body, including the liver, lungs, spleen, or brain [2, 4]. Melanocytes are responsible for the production of dark pigments on the hair, skin, eyes, and different areas of the body. Melanoma lesions tend to be black or brown in color. But in a few instances, it is also found in different colors such as azure, royal purple, rosy pink, or even colorless [5].
@@
According to the World Health Organization (WHO), cancer is one of the leading causes of death worldwide and is responsible for about 10 million deaths annually [6]. It is responsible for about 1 in every 6 deaths worldwide. Between 2008 and 2030, global cancer-related deaths are expected to go up by 45% [7]. According to the WHO, one out of every three people in the globe has skin cancer. Melanoma and non-melanoma are the most typical cancers about 300,000 new cases were found in 2018. Melanoma cost the lives of an estimated 2490 females and 4740 males in 2019[5].
@@
Deep learning (DL) is the method by which a computer model directly learns to perform categorization tasks from images, text, or sound [8, 9, 10, 11]. Multi-layered neural network architectures and labeled data are used to train models. DL lessens the requirement for some of the data preparation that machine learning generally demands [12, 13]. Unstructured data like text and images can be processed by it. Gradient descent and backpropagation are used to update and tune the DL algorithm for accuracy, enabling it to make predictions with greater precision. DL capabilities have had a significant positive impact on the healthcare sector. It has several uses in various areas of medical imaging because of its high level of precision. By using these techniques, the diagnosing procedure is expedited and human error is decreased.
@@
The most cutting-edge methods are without a doubt DL algorithms, which have lately achieved remarkable classification and segmentation results not only for medical images but also for others applications [14, 15]. So, the primary reason to create the proposed hybrid DL system is to use it as a diagnostic instrument to assist individuals who make decisions in health and medical facilities in quickly and accurately identifying melanoma in dermatological images. This reduces the burden on pathologists and hospitals at a time when the number of individuals infected with skin cancer is increasing dramatically. The following are the primary contributions of this study:
@@ 2 Literature Review
In order to detect skin cancer, Yuvika Gautam et al. [16] presented FusionEXNet, a novel fused deep-learning model. In order to take use of their capacities for feature extraction from dermoscopic images, the model combines two architectures: XceptionNet and EfficientNetV2S. Approximately 10,000 high-resolution photos from seven different kinds of skin lesions are included in the HAM10000 dataset, which was used for training and evaluation. SmoothGrad and Faster Score-CAM, two Explainable Artificial Intelligence (XAI) approaches, were integrated into the model to improve interpretability and provide insights into the model’s decision-making process. FusionEXNet demonstrated superior performance in accuracy, achieving 90.83%, compared to XceptionNet (88.82%) and EfficientNetV2S (88.01%).
@@
In order to overcome the problems of light reflections, colour fluctuations, and lesion diversity, Md Shahin Ali et al. [3] suggested a deep convolutional neural network (DCNN) model for identifying benign and malignant skin lesions from dermoscopic pictures. Using filters, preprocessing eliminates noise and artefacts from the input images, normalises them, and applies data augmentation to grow the dataset size and enhance classification accuracy. The model’s performance is contrasted with that of various other transfer learning models, such as MobileNet, AlexNet, ResNet, VGG-16, and DenseNet. The suggested DCNN model outperformed the other models in terms of robustness and reliability using the HAM10000 dataset, with testing accuracy of 91.93% and training accuracy of 93.16%.
@@
Rehan Ashraf et al. [2] suggested an intelligent Region of Interest (ROI)-based melanoma detection system to overcome the difficulties caused by nevi’s difficult-to-see visual appearance and deep learning models’ restricted data availability. The methodology ensures that only pertinent melanoma features are incorporated for system training by using an updated k-means algorithm to extract ROIs from dermoscopic pictures. This ROI-based method concentrates on the areas that include melanoma cells, which improves feature discrimination. Subsequently, a Convolutional Neural Network (CNN) transfer learning model is trained on the ROI pictures from the DermIS and DermQuest datasets using data augmentation approaches. The suggested system outperforms current techniques that depend on comprehensive picture categorisation in terms of classification accuracy.
@@
In this study [17], dermoscopy images from the HAM10000 dataset were used to create a proprietary Convolutional Neural Network (CNN) model for skin cancer classification. The seven categories bearing and malignant that the CNN model was built to identify from the photos are essential for early detection and lowering the death rate from skin cancer. An Enhanced Super Resolution Generative Adversarial Network (ESRGAN) was used to preprocess the dataset in order to improve image quality by increasing the resolution of smaller sized images.
@@
In this study [18], a novel hybrid dynamic Bayesian Deep Learning model is introduced, utilizing Three-Way Decision (TWD) theory to adaptively apply different UQ methods and neural networks in distinct classification phases. The model’s effectiveness is validated on two skin cancer datasets, achieving: accuracy of 88.95% and F1-score of 89.00% for first dataset and accuracy of 90.96% and F1-score of 91.00% for second dataset.
@@
This study [19] uses dermoscopic images to offer a two-stream deep neural network information fusion framework for multiclass skin cancer classification. A pretrained DenseNet201 model is fed with the improved image quality through a fusion based contrast enhancement technique in the first stream of the methodology, which extracts features. An approach called skewness controlled mothflame optimisation is used to optimise these properties. A refined MobileNetV2 model’s features are taken out and further downsampled in the second stream through the use of a feature selection framework. A new parallel multimax coefficient correlation technique is employed to fuse discriminant characteristics from both networks. To classify lesion images, a multiclass extreme learning machine classifier is used. Three unbalanced skin datasets HAM10000, ISBI2018, and ISIC2019, respectively are used to evaluate the concept.
@@
This article [7] introduces two novel hybrid CNN models that incorporate a support vector machine (SVM) classifier at the output layer. The approach involves: Extracting features from both CNN models. Concatenating these features and inputting them into the SVM for final classification.The proposed models outperformed existing state-of-the-art CNN models achieving accuracies of 88.02% and 87.43%.
@@
This study [20] compared the diagnostic ability of dermatologists with a Faster Region-based Convolutional Neural Network (FRCNN) trained to categorise clinical photos of pigmented skin lesions. The dataset included 3551 patients’ 5846 photos, with lesions classified into six classes: benign hematoma/hemangioma, nevus, seborrhoeic keratosis, senile lentigo, and malignant melanoma. A test set including 666 photos was generated, and the remaining 4732 images were annotated with bounding boxes to form the training set. In a six-class classification, the FRCNN model outperformed doctors with board certification (79.5%) and dermatology trainees (75.1%), with an accuracy of 86.2%. FRCNN achieved 91.5% accuracy in binary classification (benign vs. malignant), with 83.3% sensitivity and 94.5% specificity, in that order. In comparison to dermatologists, it also showed a lower false positive rate of 5.5%. The results show that the FRCNN model performed better in picture classification tasks than 20 dermatologists, indicating that the public may benefit from using it in the future to improve skin cancer prognosis.
@@
The goal of this work [21] is to create deep learning models for the classification of dermal cell pictures in order to detect skin cancer. Deep learning techniques are used at the heart of the system to improve prediction accuracy through the use of a model-driven architecture hosted in the cloud. Building these models and using them to categorise dermal cells constitute the process; standard datasets are used for testing. With this method, practitioners may quickly develop and use deep learning models for early skin cancer identification.
@@
This work [4] addressed the difficulties caused by the fine-grained heterogeneity in diagnostic categories by developing a preprocessing pipeline for multiclass skin cancer classification. Image hair removal, dataset enhancement, and image resizing to accommodate various models were all part of the pipeline. To fine-tune the EfficientNet B0-B7 models for classification, transfer learning was used on pre-trained ImageNet weights using the HAM10000 dataset. With an F1 Score of 87% and Top-1 Accuracy of 87.91%, EfficientNet B4 had the best results when the models’ performance was assessed using Precision, Recall, Accuracy, F1 Score, and Confusion Matrices. More complicated models do not always perform better; models with intermediate complexity, such EfficientNet B4 and B5, showed notable superiority over more complex models.
@@
This research [22] utilizes images from various publicly available ISIC datasets to create a balanced dataset consisting of 10,500 images for training and testing. An ensemble of four convolutional neural network (CNN) architectures-ResNet50, EfficientNet B6, InceptionV3, and Xception was employed for melanoma classification.The experimental results demonstrate that the proposed ensemble model effectively classifies melanoma skin cancer with high accuracy, outperforming many state-of-the-art methods.
@@ 3.1 Dataset
Our dataset basically consists of two distinct categories of skin carcinoma images. DL requires a significant amount of data in order to produce a good result. The collection of skin cancer pictures, on the other hand, is critical. To address these issues, we used a freely available dataset of 3297 dermoscopic images from Kaggle [23]. It includes 1800 images of benign and 1497 images of malignant categorized moles. The data was obtained via ISIC Archive. Each picture is 224x244 pixels in size. The professional pathologist confirmed the database’s ground truth.
@@ 3.2 Image Preprocessing
The main goal of the image preprocessing process is to make the data suitable for deep learning models and improve the initial medical images by getting rid of noise, air bubbles, and artifacts caused by gel that was sprayed before the picture was taken. To attain a high classifying rate, we removed artifacts and noise from the pictures. The elimination of this noise and artefacts was crucial to guaranteeing high-quality input images since they could mask important features required for classification. The image was smoothed using noise reduction algorithms, which also removed any random deviations that could confound the model. We employed feature scaling and turned the categorical label data into number data[15]. In order to ensure that the input data had consistent numerical ranges, we used feature scaling to normalise the pixel intensity values, which can vary significantly in photos. This step prevents gradient instability problems and helps models converge more quickly during training.
@@ 3.2.1 Class Labeling
We used label encoding techniques to transform the categorical data into numerical format because the labels in medical picture datasets are typically categorical (e.g., different forms of skin cancer). In order to enable the deep learning model to process and learn from the labels during training, each distinct category was given a numerical value. The labeling data collection is presented as follows: the image’s label 0 indicates benign patients, while 1 symbolizes malignant patients[24].
@@ 3.2.2 Feature scaling
The feature scaling method is well recognized in the field of machine learning and pattern recognition for being used to normalize data. To avoid outliers, this approach sets all data items to the same scale and thus improves prediction quality. As the features in cancer datasets have a high variance, One of the pre-processing strategies for normalization is feature scaling. We divided the grayscale value of an image by 255 to standardize our image pixels between 0 and 1. As a result, the numbers will be relatively little, and the computation will be simpler and faster[25, 26].
@@ 3.3 Transfer Learning
Transfer Learning has emerged as a popular method in Deep Learning (DL) that utilizes a pre-trained model for a similar task to train a new model. This approach is particularly useful in medical imaging, where training on large datasets such as ImageNet with all neural network parameters can be computationally intensive [6]. The ImageNet dataset, for instance, contains 14 million images with 20,000 categories for visual recognition tasks [27].
In this study, we propose a model for skin lesion classification, which was trained and evaluated using advanced Convolutional Neural Networks (CNNs) such as Inception V3, Densenet121, and a Hybrid model [21, 3]. To optimize our model, we replaced the last softmax layer with a sigmoid activation, reducing the classification output to two classes for identifying benign and malignant skin lesions.
The Inception V3 CNN model achieved remarkable accuracy of over 78.1% on the ImageNet dataset and has been widely used for image classification tasks [25]. Its architecture includes inception modules that use filters of different sizes on the same input, and auxiliary classifiers for regularization weight loss ratio. The convolution layer implements factorization to minimize dimensionality and avoid overfitting.
Another CNN architecture utilized in our model is DenseNet121, which resolves the vanishing gradient problem by directly connecting every layer to every other layer. Instead of summing feature maps from previous layers, DenseNet121 concatenates them, resulting in a compact model with fewer parameters and enabling feature reuse [28].
@@ 3.4.1 Inception V3
InceptionV3 is a CNN-based DL model used for image classification. It is a highly renowned image detection model that has been developed over time by a number of researchers. On the ImageNet dataset, the model achieved more than 78.1% accuracy. Google’s InceptionV3 is the third version in the DL Convolutional Architecture series. The model was developed based on 1000 classes from the ImageNet dataset, which was trained on approximately 1 million images. It is made up of inception modules that apply filters of various sizes to the same amount of input. In the Inception network, the auxiliary classifiers add to the regularization weight loss ratio. Factorization was implemented in the convolution layer in order to minimize dimensionality and therefore overfitting [25].